OpenStudy (anonymous):

Solve the compound inequality 60 < 20u < 120. Question 1 options: 1) 3 < u < 6 2) 6 > u > 7 3) 6 < u < 3 4) 3 > u < 6

OpenStudy (ash2326):

We have \[60<20u<120\] Divide all sides by 20 \[\frac{60}{20}<\frac{20u}{20}<\frac{120}{20}\] @RUTHIEBOO18 Can you solve now?